UWU Company was incorporation on January 1, 2021. The following were the transactions during the year:

- Total consideration from share issuance amounted to P4,000,000.
- A land and building were acquired through a lump sum payment of P800,000. A mortgage amounting to P200,000 was assumed on the land and building.
- Total payments of P160,000 were made during the year on the mortgage assumed n the land and building. The payments are inclusive of interest amounting to P20,000.
- Additional capital of P400,000 was obtained through bank loans. None of the bank loans were paid during the year. Half of the bank loan required a secondary mortgage on the land and building.
- There is no accrued interest as of year-end.
Dividends declared during the year but remained unpaid amounted to P120,000.
- No other transactions during the year affected liabilities.
- Retained earnings as of December 31, 2021 is P240,000

How much is the total assets as of December 31, 2021?
Answer #1

Answer

Total assets=Total liabilities +Shareholder's equity

Shareholder equity=Common stock+Retained earnings

=P 4,000,000+240,000

Shareholder equity=P 4,240,000

Total liabilities=Mortgage on asset atken-Loan Repaid+Bank loan taken+Dividend payable

=200,000-140,000+400,000+120,000

Total liabilities=P 580,000

Note: Amount of laon paid of160,000 includes 20,000 so 20,000 is deducted before calculating liability amount apdi

Loan repaid=160,000-20,000=140,000

Total assets=P 580,000+P 4,240,000

Total assets=P 4,820,000